A stone is thrown vertically upward from the top of the building. If the equation of the motion of the stone is 𝑠(𝑡) = −5𝑡2 + 30𝑡 + 200 , where 𝑠 is the directed distance from the ground in meters and 𝑡 is in seconds.


a. What is the height of the building?

b. What is the average velocity on the time interval [1,3]

c. What is the instantaneous velocity at time t=1?

d. at what time the stone will hit the ground?

e. What is the instantaneous velocity of the stone upon impact?


1
Expert's answer
2021-08-20T12:56:53-0400

a. At "t = 0" the value of "s(t)" gives the height of the building. Thus:


"h = s(0) = 200m"

b. The distance from the ground at "t = 1" is:


"s(1) = -5\cdot 1^2 + 30\cdot 1 + 200 = 225m"

The distance from the ground at "t = 3" is:


"s(3) = -5\cdot 3^2 + 30\cdot 3 + 200 = 245m"

The average velocity is:


"v_{av} = \dfrac{s(3) - s(1)}{3s - 1s} \approx 6.7m/s"

c. The instantaneous velocity is:


"v(t) = s'(t) = -10t+30\\
v(1) = -10+30 = 20m/s"

d. The stone will hit the ground when "s(t) = 0." Solving the following quadratic equation for "t", obtain:


"-5t^2 + 30t + 200 = 0\\
t = 10s"

c. The instantaneous velocity at this time is:


"v(10) = -10\cdot 10 + 30 = -70m/s"